# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v1

@@ 7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two clusters (k=2), with silhouette scores tightly concentrated between 0.82 and 0.83 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets. In contrast to the Breast Cancer dataset, the normalized entropy is consistently lower (approximately 0.56-0.58), indicating that explanation modes are unequally populated rather than evenly balanced.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.67 to 0.69, with standard deviations on the order of $10^{-3}$ or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig.3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
[…]
@@
As a control, fixing the regularisation strength results in a collapse to a single explanation cluster across all splits (refer to the results folder in the provided code repository). Therefore, the COMPAS results show that explanation multiplicity in Logistic Regression is highly consistent across splits and strongly tied to distinct regularisation regimes, while predictive performance remains effectively unchanged.
Figure 3: COMPAS — Logistic Regression (single split). Left: PCA embedding of explanation (SHAP) vectors obtained from repeated training runs on a fixed train–test split (split 105). Points are coloured by explanation cluster assignment. Red stars indicate representative cluster centroids, and the black $\times$ denotes the average attribution vector across all runs. Right: The same embedding with predictive accuracy overlaid. Accuracy remains within a narrow range across the explanation manifold, indicating that distinct explanatory modes coexist at similar predictive performance.
Figure 4: COMPAS - Logistic Regression: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking model runs across all train-test splits. Each point corresponds to a trained model instance and is coloured by predictive accuracy. Multiple separated explanation tracks persist across splits, indicating that explanation variability is not confined to a single resampling of the data. The black $\times$ marks the global average attribution vector, computed as the arithmetic mean of all SHAP vectors across runs and clusters; this point lies away from the populated tracks of the manifold.
@@
Table 3: Split-level summary for Logistic Regression on COMPAS under stochastic training. For each train-test split, we report the number of runs, the selected number of explanation clusters $k^{*}$, the silhouette score at $k^{*}$, the normalized cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). This table quantifies split-to-split consistency of the clustering outcome and confirms that multiplicity is not driven by unstable performance.
Split Runs $k^{*}$ Silh. $(k^{*})$ $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 2 0.829 0.568 0.683 0.001
101 1000 2 0.830 0.568 0.685 0.001
102 1000 2 0.830 0.566 0.694 0.001
103 1000 2 0.822 0.574 0.671 0.001
104 1000 2 0.826 0.566 0.681 0.001
105 1000 2 0.824 0.579 0.679 0.001
106 1000 2 0.830 0.566 0.676 0.000
107 1000 2 0.830 0.568 0.678 0.003
108 1000 2 0.830 0.574 0.674 0.002
109 1000 2 0.829 0.571 0.686 0.001
@@
100 0 866 12.578 21.767 [0.034, 100.0] 0.683 0.000
100 1 134 0.020 0.007 [0.010, 0.0] 0.682 0.001
101 0 866 12.578 21.767 [0.034, 100.0] 0.685 0.000
101 1 134 0.020 0.007 [0.010, 0.0] 0.684 0.002
102 0 867 12.564 21.759 [0.034, 100.0] 0.694 0.000
102 1 133 0.020 0.007 [0.010, 0.0] 0.692 0.001
103 0 864 12.607 21.784 [0.035, 100.0] 0.671 0.001
103 1 136 0.020 0.007 [0.010, 0.0] 0.671 0.001
104 0 867 12.564 21.759 [0.034, 100.0] 0.682 0.000
104 1 133 0.020 0.007 [0.010, 0.0] 0.681 0.002
105 0 862 12.637 21.801 [0.036, 100.0] 0.679 0.001
105 1 138 0.020 0.007 [0.010, 0.0] 0.676 0.001
106 0 867 12.564 21.759 [0.034, 100.0] 0.676 0.000
106 1 133 0.020 0.007 [0.010, 0.0] 0.676 0.001
107 0 866 12.578 21.767 [0.034, 100.0] 0.679 0.001
107 1 134 0.020 0.007 [0.010, 0.0] 0.672 0.003
108 0 864 12.607 21.784 [0.035, 100.0] 0.674 0.001
108 1 136 0.020 0.007 [0.010, 0.0] 0.670 0.002
109 0 865 12.593 21.776 [0.035, 100.0] 0.687 0.001
109 1 135 0.020 0.007 [0.010, 0.0] 0.683 0.001
7.1.3 Disagreement between equally accurate basins
@@
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each held-out test instance $i$, we compute the absolute difference in predicted class probability,
$\Delta_{i}=\left|P(y=1\mid x_{i},M_{b_{1}})-P(y=1\mid x_{i},M_{b_{2}})\right|,$ (1)
where $M_{b_{1}}$ and $M_{b_{2}}$ denote representative models from two different explanation clusters.
@@
On the Breast Cancer dataset, disagreement magnitudes span a wide range. Across held-out test instances, the mean disagreement is modest, but the upper tail is substantial. In split 104, the maximum disagreement reaches $\Delta=0.85$, with several cases exceeding $\Delta=0.4$. These large discrepancies occur despite both models achieving similarly high overall accuracy. The highest-disagreement cases correspond to specific individuals for whom one model assigns a near-certain malignant (or benign) probability while another assigns a substantially lower risk, indicating that explanation multiplicity can lead to sharply divergent predictions at the individual level.
On the COMPAS dataset, disagreement magnitudes are smaller but non-zero. For split 105, the mean disagreement across test instances is approximately 0.01, with a maximum disagreement of 0.05. Although these values are lower than those observed for Breast Cancer, they still indicate that equally accurate models can assign different recidivism probabilities to the same individual, even when explanation clusters are strongly separated and predictive accuracy is tightly constrained.
@@
These results show that explanation multiplicity is not merely a geometric property of explanation space, but can translate into different model outputs for individual test instances. In settings where multiple explanation basins coexist at similar accuracy, the choice of model instance implicitly determines which individuals receive higher or lower predicted risk, even when aggregate performance metrics are indistinguishable. These clusters therefore represent qualitatively distinct explanatory mechanisms rather than perturbations of a single solution. Despite near-identical test accuracy, the two basins induce materially different prediction behaviour at the individual level, with a maximum disagreement of 0.848. This demonstrates that high predictive performance alone does not imply mechanistic stability or uniqueness.
[…]
@@
To understand the source of these disagreements, we examine the characteristic explanation patterns associated with each basin. Figure 5 visualises the centroid attribution profiles of the two explanatory modes identified on the Breast Cancer dataset, revealing how equally accurate models rely on qualitatively different feature groupings when making predictions.
@@ 7.2 Random Forests: explanation multiplicity in ensemble models
Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class. Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling. Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more complex and widely deployed learning architectures.
@@ 7.2.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Table 5; Fig. 7 illustrates that the same multi-modal structure is already present within a single train-test split. Each train-test split aggregates approximately 192,000 Random Forest instances, reflecting stochasticity arising from bootstrap sampling, feature subsampling, and hyperparameter variation. In explanation space, SHAP vectors form multiple dense and well-separated regions rather than the smooth manifolds observed for Logistic Regression. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting.
Table 5 shows that two explanation clusters are consistently identified in each split, with a stable size ratio of approximately 1:2. Predictive performance remains comparable across clusters: mean accuracy differences are small (typically below 0.01), and standard deviations overlap substantially. The global average attribution vector lies outside the high-density regions of the explanation manifold (Fig. 6), indicating that averaging explanations across stochastic runs again produces a representation that does not correspond to a typical deployed model instance. Together, these results show that explanation multiplicity persists in Random Forests despite ensemble averaging and high predictive accuracy.
@@
Figure 6: Breast Cancer - Random Forest: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking Random Forest models across all train–test splits. Each point corresponds to a trained ensemble instance and is coloured by predictive accuracy. Multiple dense and separated regions are visible in explanation space, indicating persistent explanation multiplicity across resampled datasets. The black cross denotes the global average attribution vector, computed as the mean of all SHAP vectors across runs, which lies outside the high-density regions of the manifold.
Table 5: Cluster-level summary for Random Forest on the Breast Cancer dataset under stochastic training with varied hyperparameters. Each row corresponds to one explanation cluster within a train-test split. We report cluster size and predictive accuracy (mean $\pm$ std across runs assigned to the cluster). Across splits, two explanation clusters are consistently observed, with comparable predictive performance and stable relative sizes.
Split Cl. Runs Acc. mean Acc. std
100 0 64168 0.967 0.005
100 1 127832 0.974 0.005
101 0 64104 0.956 0.008
101 1 127896 0.959 0.006
102 0 64031 0.953 0.011
102 1 127969 0.956 0.005
103 0 64120 0.946 0.012
103 1 127880 0.949 0.007
104 0 64145 0.941 0.007
104 1 127855 0.945 0.007
Figure 7: Random Forest explanation landscape for Breast Cancer (single split). PCA projection of explanation vectors for split 103 shows two distinct, high-density regions within a single train–test split. The first principal component explains 75.8% of the variance, while PC2 captures residual spread. Colour indicates test accuracy, which varies smoothly across both regions. The persistence of multiple explanation clusters within a single split, and at comparable accuracy, indicates that mechanistic multiplicity in Random Forests is intrinsic and does not rely on aggregation across splits or hyperparameter overlay.

### v2

@@ 7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two clusters (k=2k{=}2), with silhouette scores ranging from 0.73 to 0.91 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets—notably higher than those observed on Breast Cancer. The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with cluster sizes typically in a ratio of approximately 1:2.5.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.68 to 0.70 (this is typical for recidivism prediction tasks given inherent unpredictability), with standard deviations on the order of $10^{-3}$ or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig. 3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
Table 4 shows that, as with Breast Cancer, both explanation clusters span the full range of regularisation strengths. In every split, both clusters contain runs with $C$ values ranging from 0.01 to approximately 100, with similar mean $C$ values of around 10–12. This indicates that the mechanistic separation in explanation space is not a simple partition by regularisation strength. Rather, the two explanation modes represent genuinely distinct regions of the explanation landscape that models can reach regardless of their regularisation setting. Predictive accuracy remains effectively identical across clusters within each split, with differences typically below 0.001.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength ($C=1.0$). Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster ($k=1$), with zero entropy and negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or clustering procedure, but emerges specifically when the training process admits multiple equally accurate models through varied regularisation.
Therefore, the COMPAS results reinforce the findings from Breast Cancer: explanation multiplicity in Logistic Regression is highly consistent across splits, while predictive performance remains effectively unchanged. Importantly, the mechanistic separation is not driven by hyperparameter differences—models with similar $C$ values can land in different explanation basins, indicating that multiplicity reflects the structure of the explanation landscape itself.
Figure 3: COMPAS - Logistic Regression (split 109). Left: PCA embedding of explanation (SHAP) vectors obtained from 1,000 independent training runs on a fixed train-test split. Points are coloured by cluster assignment identified in normalised explanation space. Silhouette-based selection yields $k=2$ clusters with silhouette score 0.727 and normalised entropy $H_{\mathrm{norm}}=0.844$, indicating well-separated but unequally populated explanation modes. Cluster 0 (blue, $n=272$, mean $C=10.56$) and Cluster 1 (orange, $n=728$, mean $C=11.58$) occupy distinct regions of the explanation manifold, forming a curved arc geometry similar to that observed on Breast Cancer. Stars indicate cluster centroids; the black cross denotes the average attribution vector across all runs, which lies between the two clusters in a sparsely populated region. Right: The same embedding with regularisation strength $C$ overlaid. Both clusters span the full range of $C$ values, with a smooth gradient visible along the arc, indicating that the mechanistic separation reflects the structure of the explanation landscape rather than a simple partition by regularisation strength. The first two principal components capture 79.3% and 15.9% of variance respectively (95.2% total), providing a faithful low-dimensional representation of the explanation geometry.
Figure 4: COMPAS — Logistic Regression: universal explanation manifold. PCA embedding of SHAP vectors obtained by stacking N=10, [^0] N=10{,}000 training runs across 10 independent train–test splits. Top-left: Points coloured by split seed show distinct trajectories in explanation space for each data realisation. Unlike the Breast Cancer dataset, COMPAS exhibits greater between-split variability in manifold geometry, reflecting sensitivity to the particular train–test partition. Top-right: Accuracy overlay shows predictive performance confined to a narrow range (0.67–0.70) across the entire manifold, confirming that distinct explanatory modes coexist at comparable accuracy. Bottom-left: Raw regularisation strength $C$ is distributed across the manifold without clear spatial structure within splits. Bottom-right: $\log_{10}(C)$ overlay shows that both strong regularisation (blue, $\log_{10}C<0$) and weak regularisation (red, $\log_{10}C>0$) appear throughout the explanation space. The first two principal components capture 47.4% and 24.9% of variance respectively, indicating a more complex explanation geometry than observed for Breast Cancer. The persistence of multiple explanation tracks across splits confirms that mechanistic multiplicity on COMPAS is a stable property of the model class.
@@
Table 3: Split-level summary for Logistic Regression on COMPAS under stochastic training with within-split clustering. For each train–test split, we report the number of runs, the selected number of explanation clusters $k$, the silhouette score at $k$, the normalised cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). All splits consistently yield $k=2$ explanation clusters with high silhouette scores (0.73–0.91), indicating stable and well-separated mechanistic multiplicity. Predictive accuracy remains tightly constrained (0.68–0.70) across all splits and clusters.
Split Runs $k$ Silh. ($k$) $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 2 0.802 0.834 0.693 0.001
101 1000 2 0.910 0.841 0.695 0.002
102 1000 2 0.743 0.908 0.691 0.001
103 1000 2 0.814 0.846 0.684 0.001
104 1000 2 0.803 0.870 0.689 0.003
105 1000 2 0.833 0.897 0.688 0.001
106 1000 2 0.767 0.897 0.682 0.001
107 1000 2 0.873 0.855 0.685 0.001
108 1000 2 0.841 0.855 0.684 0.002
109 1000 2 0.727 0.844 0.698 0.001
@@
100 0 735 9.338 18.396 [0.010, 99.4] 0.693 0.001
100 1 265 12.323 22.541 [0.010, 99.2] 0.693 0.001
101 0 730 12.290 21.532 [0.010, 98.8] 0.695 0.001
101 1 270 9.964 20.581 [0.010, 99.9] 0.695 0.002
102 0 677 11.069 20.082 [0.010, 99.6] 0.691 0.001
102 1 323 12.180 22.227 [0.010, 99.5] 0.691 0.001
103 0 727 11.190 20.780 [0.010, 99.0] 0.684 0.001
103 1 273 10.901 21.966 [0.010, 99.3] 0.684 0.001
104 0 709 11.658 21.687 [0.010, 99.8] 0.689 0.003
104 1 291 11.129 20.762 [0.010, 92.7] 0.689 0.003
105 0 687 11.334 20.312 [0.010, 97.1] 0.688 0.001
105 1 313 10.141 20.824 [0.010, 96.9] 0.688 0.001
106 0 687 10.592 21.029 [0.010, 98.9] 0.682 0.001
106 1 313 10.476 19.777 [0.010, 97.4] 0.682 0.001
107 0 720 11.347 20.892 [0.011, 97.2] 0.685 0.001
107 1 280 10.400 18.852 [0.010, 89.9] 0.685 0.001
108 0 720 11.047 21.038 [0.010, 99.7] 0.684 0.002
108 1 280 8.642 18.258 [0.011, 93.3] 0.685 0.002
109 0 728 11.576 20.999 [0.010, 98.3] 0.698 0.001
109 1 272 10.562 20.295 [0.012, 99.7] 0.698 0.001
7.2 Disagreement Between Equally Accurate Explanation Basins
@@
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each held-out test instance $i$, we compute the absolute difference in predicted malignant probability,
$\Delta_{i}=\left|P(y=1\mid x_{i},M_{1})-P(y=1\mid x_{i},M_{2})\right|,$
where $M_{1}$ and $M_{2}$ denote representative models selected from two different explanation clusters. Importantly, the compared models share the same model class, training data, and optimisation objective, and achieve comparable test accuracy. This controls for performance differences and isolates disagreement attributable to distinct explanatory mechanisms rather than optimisation noise or underfitting.
@@
On the Breast Cancer dataset, disagreement magnitudes span a wide range. While the mean disagreement across test instances is typically modest, the upper tail is substantial. Figure 5 illustrates this phenomenon for a representative split (Split 107), in which 76 test instances exhibit disagreement exceeding $\Delta>0.1$, 33 exceed $\Delta>0.2$, and 15 exceed $\Delta>0.3$. The maximum observed disagreement in this split reaches $\Delta=0.68$, despite both representative models achieving similar overall accuracy (approximately 0.95). These values indicate that explanation multiplicity can manifest as large probability divergences for individual patients even when aggregate performance metrics are indistinguishable.
The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite clinical decisions for the same patient. Notably, this disagreement does not arise from marginal predictions near the decision threshold: one model is confidently malignant, while the other is confidently benign. The accompanying SHAP attributions reveal qualitatively different patterns of feature reliance across the two explanation basins. Although both models attend to overlapping morphological feature families, they weight and combine this evidence differently, resulting in incompatible diagnostic conclusions. This behaviour is inconsistent with an interpretation of disagreement as stochastic fluctuation or local instability, and instead reflects the presence of distinct explanatory mechanisms that remain viable at comparable accuracy.
Across all ten train–test splits, the magnitude of individual-level disagreement varies substantially. Maximum disagreement per split ranges from negligible values (e.g., $\Delta\approx 0.05$ in split 103) to large discrepancies exceeding $\Delta=0.7$ (e.g., splits 101 and 107). This variability indicates that the practical manifestation of explanation multiplicity depends on the particular data realisation. However, critically, large individual-level disagreements can arise even when aggregate accuracy and calibration metrics remain effectively unchanged. In such cases, standard performance summaries obscure the fact that different model instances may assign substantially different risk estimates to the same individual.
Crucially, disagreement is not confined to cases where models predict different labels. Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities ($\Delta\approx 0.21$). In these cases, predictions agree at the label level, but differ meaningfully in confidence, with one model indicating borderline risk and the other indicating high certainty. While these instances do not lead to immediate label flips, they demonstrate that mechanistic disagreement can persist even when classification outcomes are identical, manifesting as divergent belief states rather than discrete decision changes. Such differences are relevant in downstream clinical contexts where predicted probabilities inform risk stratification, follow-up decisions, or human oversight.
@@
Taken together, these results show that explanation multiplicity is not merely a geometric or cosmetic property of explanation space. When multiple explanation basins coexist at similar accuracy, the choice of model instance can materially affect individual-level predictions, either by altering the predicted class or by substantially shifting estimated risk. The observed clusters therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution. High predictive performance alone does not imply mechanistic stability or explanatory uniqueness.
@@
To further characterise the structure of these mechanisms, we examine the characteristic explanation patterns associated with each basin. Figure 5 visualises the centroid attribution profiles of the explanatory modes identified on the Breast Cancer dataset, highlighting how equally accurate models systematically rely on different feature groupings when forming predictions.
Figure 5: Mechanistic disagreement between equally accurate explanation basins (Breast Cancer, Split 107). Disagreement heatmap (left) showing predicted malignant probabilities for two explanation clusters ($C_{0}$, $C_{1}$), sorted by absolute disagreement $\Delta$. The highlighted case illustrates a large disagreement ($\Delta=0.677$) for a malignant test instance, where one cluster confidently predicts malignancy ($P=0.72$) while the other confidently predicts benign ($P=0.04$), resulting in opposite clinical decisions despite comparable overall accuracy. Right panels show the corresponding mean absolute SHAP attributions for representative models from each cluster, revealing distinct feature attribution profiles underlying the disagreement. This example demonstrates that mechanistic multiplicity can yield materially different predictions for the same patient, and that such disagreement is not confined to marginal probability regimes.
Table 5: Same-label probability disagreement among correctly classified test samples (Breast Cancer, Split 107): We show two representative instances to illustrate that two equally accurate explanation basins can assign the same predicted class and correctly classify the outcome, yet differ substantially in predicted malignant probability ($\Delta=0.21$). These examples demonstrate that mechanistic disagreement persists even when predictions and accuracy are stable, appearing as materially different confidence assessments rather than label flips.
Test index True label $C_{0}$ $P(\text{malig})$ $C_{1}$ $P(\text{malig})$ $\Delta$
36 Benign 0.79 1.00 0.21
169 Malignant 0.60 0.81 0.21
@@ 7.3 Random Forests: explanation multiplicity in ensemble models
Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class. Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling. Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more expressive and widely deployed ensemble architectures.
@@ 7.3.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Fig. 7. Each train–test split aggregates approximately $2\times 10^{5}$ Random Forest instances, reflecting stochasticity arising from bootstrap resampling, feature subsampling, and hyperparameter variation. This large-scale sampling allows the structure of explanation space to be examined at high resolution.
Figure 6 visualises the universal explanation manifold obtained by stacking explanation vectors across multiple train–test splits. In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into multiple dense and well-separated regions. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting. As in the linear case, the global average attribution vector—computed by averaging SHAP values across all runs—lies outside the high-density regions of the manifold. This indicates that aggregating explanations across stochastic training runs again produces a representation that does not correspond to a typical deployed model instance.
@@
Figure 6: Breast Cancer — Random Forest: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking Random Forest models across all train–test splits under stochastic training. Each point corresponds to one trained ensemble instance and is coloured by test accuracy. Explanation vectors organise into multiple dense and well-separated regions that persist across splits, indicating stable explanation multiplicity in an ensemble model class. Predictive accuracy varies smoothly across the manifold and does not align with any single region. The black cross denotes the global average attribution vector, computed by averaging SHAP values across all runs, which lies outside the high-density regions of the explanation space.
Importantly, predictive accuracy varies smoothly across the explanation manifold and does not align with any single cluster or region. Models drawn from different explanation clusters achieve comparable test accuracy, with no sharp performance boundary separating the modes. This demonstrates that, in Random Forests, explanation multiplicity is not driven by accuracy trade-offs or poorly performing solutions, but instead reflects the coexistence of distinct explanatory mechanisms within the same performance regime.
Figure 7 shows the explanation landscape for a single representative split (Split 104). Even within a fixed train–test realisation, explanation vectors again separate into two high-density regions. The first principal component captures the majority of variance in explanation space, while the second component reflects residual spread within each mode. As in the universal manifold, predictive accuracy varies continuously across both regions, indicating that explanation multiplicity arises intrinsically within a single split and does not depend on aggregation across data resamplings or extreme hyperparameter configurations.
Taken together, these results show that explanation multiplicity persists in Random Forests despite ensemble averaging and strong inductive bias towards prediction stability. In contrast to Logistic Regression, Random Forests often exhibit limited disagreement in predicted labels or probabilities at the individual level, reflecting the stabilising effect of ensembling. However, this output stability masks substantial mechanistic non-uniqueness: multiple, distinct explanation basins coexist at comparable accuracy. Thus, even when predictions appear stable, the underlying reasoning employed by the model remains non-unique.
Figure 7: Breast Cancer — Random Forest: explanation landscape within a single train–test split. PCA projection of explanation (SHAP) vectors for Random Forest models trained on a single representative split (Split 104). Two high-density regions are visible within the explanation space, demonstrating that explanation multiplicity arises intrinsically within a fixed data realisation and does not rely on aggregation across splits. Colour indicates test accuracy, which varies continuously across both regions without sharp separation. The global average attribution vector (black cross) again lies outside the dense regions, indicating that explanation averaging does not correspond to a typical model instance even within a single split.
This contrast highlights an important distinction between predictive stability and mechanistic stability. Ensemble methods may suppress variability in outputs while preserving—or even amplifying—multiplicity in explanation space. High predictive accuracy and low output variance therefore do not imply a unique or well-defined explanatory mechanism.